OpenStudy (anonymous):

use substitution x=2tan(theta) to solve the integral dx/((x^2(x^2+4)^(1/2)) from 1/2 to 1

OpenStudy (anonymous):

\[\large \int_{1}^{\frac {1}{2}}\frac {dx}{x^2\sqrt{x^2+4}} \] that's the integral?

OpenStudy (anonymous):

 can you simplify and do it from here?